DUALIZING DG MODULES AND GORENSTEIN DG ALGEBRAS

ANDERS FRANKILD, SRIKANTH IYENGAR, AND PETER JØRGENSEN

Introduction

There is a growing body of research concerned with extending the rich theory of commutative Gorenstein rings to DG (=Differential Graded) algebras. This subject began with the work of Félix, Halperin, and Thomas [7] on a Gorenstein condition for (cochains complexes of) topological spaces, which extends classical Poincaré duality for manifolds. Their study was complemented by that of Avramov and Foxby [3], who adopted a similar definition of the Gorenstein property, but focused on DG algebras arising in commutative ring theory. Further progress along these lines has been made by Dwyer, Greenlees, and Iyengar [6].

In a recent article, Frankild and Jørgensen [13] proposed a new notion of ‘Gorenstein’ DG algebras. The natural problem arises: How does their approach relate to those of Félix-Halperin-Thomas and Avramov-Foxby? The content of one of the main theorems in this paper is that the Avramov and Foxby definition of Gorenstein, for the class of DG algebras considered by them, is equivalent to the one of Frankild-Jørgensen.

In order to facilitate the ensuing discussion, we adopt the convention that the term ‘Gorenstein DG algebra’ is used in the sense of Frankild and Jørgensen. Their definition is recalled in Section 4; see also the discussion below. Theorem I. Let R = {R_{i}}_{i ⩾ 0} be a commutative DG algebra where R_{0} is a local ring, with residue field k, and the H_{0}(R)-module H(R) is finitely generated.

Then R is Gorenstein if and only if rank_{k} Ext_{R} (k, R) = 1.

The implication: when R is Gorenstein the k-vector space Ext_{R} (k, R) is one dimensional, was proved in [13]. The converse settles a conjecture in loc. cit.

We establish a similar result for cochain complexes of certain topological spaces. This amounts to the statement: Let k be a field and let X be a finite simply connected space. If X is Gorenstein at k, in the sense of Félix, Halperin, and Thomas, then the cochain complex of X (with coefficients in k) is Gorenstein. The converse holds if the characteristic of k is 0.

Theorem I is proved in Section 4, as is a corollary related to ascent and descent of the Gorenstein property. The topological counterpart of Theorem I is to be found in Section 5.

By definition, a DG algebra is Gorenstein if it is dualizing, in the sense of [11], as a module over itself. Thus, it is no surprise that the protagonists in this work are dualizing DG modules over DG algebras. These objects, which are DG analogues of dualizing complexes over rings (which may be viewed as DG algebras concentrated in degree 0), were introduced in loc. cit., wherein one finds also DG versions of some central results in the theory of dualizing complexes for rings; see also [10], [12].

Date: December 11, 2002.

2000 Mathematics Subject Classification. 13D99,13H10,16E45.

Key words and phrases. dualizing DG modules, Gorenstein DGA, Gorenstein space, ascentdescent of Gorenstein property for DGAs.

Part of this research was done when S. I. was at the University of Sheffield, supported by a grant from the E. P. S. R. C.

Section 3 delves deeper into this subject, focusing on finite commutative local DG algebras: a class of DG algebras which is slightly larger than that considered in Theorem I. For instance, one finds there the following result that detects dualizing DG bimodules. The adjective ‘symmetric’ is explained in (1.1).

Theorem II. Let R be a finite commutative local DG algebra and D a symmetric DG R-bimodule. Then D is dualizing if and only if the H_{0}(R)-module H(D) is finitely generated and rank_{k} Ext_{R} (k, D) = 1.

This is one of the principal results of this section, and indeed of this work. One piece of evidence for its efficacy is that one can immediately deduce Theorem I from it. Section 3 contains also the following result:

Theorem III. Let R be a finite commutative local DG algebra. Then any pair of symmetric dualizing DG R-bimodules are quasiisomorphic up to suspension.

The results obtained demonstrate that the theory of dualizing DG modules over finite commutative local DG algebras is akin to that over commutative local rings.

There are numerous questions that have not been addressed as yet, especially when one considers dualizing DG modules over general, not necessarily commutative, DG algebras. Among these, perhaps the most important one is: Do dualizing DG modules exist?

Section 2 deals exclusively with this question. The following portmanteau result sums up the results arrived at there; it shows that many of the naturally arising DG algebras possess dualizing DG modules. The paper is so organized that statements (1)–(4) are in Section 2, while the last one is in Section 5.

Theorem IV. Let A be a commutative noetherian ring with a dualizing complex. The following statements hold:

(1) The Koszul complex on each finite set of elements in A admits a dualizing DG module;

(2) For each bounded complex of finitely generated projective A-modules P with H(P) ≠ 0, the endomorphism DG algebra Hom_{A} (P, P) admits a dualizing DG module;

(3) Suppose that A is local and that φ: A′ → A is a local homomorphism of finite flat dimension. The DG fibre of φ admits a dualizing DG module;

(4) The chain complex (with coefficients in A) of each finite topological monoid admits a dualizing DG module;

(5) The cochain DG algebra (with coefficients in a field) of each finite simply connected CW complex admits a dualizing DG module.

Besides being of intrinsic interest, we expect the existence result to be a useful tool in DG homological algebra. Already, (5) is critical for the work in Section 5.

As is to be expected, extensive use is made of techniques from homological algebra. For ease of reference, Section 1 recapitulates some of the notions and constructions in this subject that arise frequently in this paper.

1. DG homological algebra

It is assumed that the reader is familiar with basic definitions concerning DG algebras and DG modules; if this is not the case, then they may consult, for instance, [8]. Moreover, in what follows, a few well known results in this subject are used; for these we quote from [4] whenever possible, in the interest of uniformity.

A few words about notation: Most graded objects that appear here are assumed to be graded homologically; for instance, a graded set X is a collection of sets {X_{i}}_{i∈ℤ}. Also, in a complex M, the differential ∂ decreases degree: ∂_{i}: M_{i} → M_{i−1}. All exceptions to this convention are to be found in Section 5 that deals with examples from topology, where the natural grading is the cohomological one.

1.1. Modules. Over a DG algebra R it is convenient to refer to left DG Rmodules (respectively, right DG R-modules) as left R-modules (respectively, right R-modules), thus omitting the ‘DG’. One exception: When speaking of a ring, rather than a DG algebra, we distinguish between modules, in the classical sense of the word, and complexes, which are the DG modules.

We denote R^{∘} the opposite DG algebra. Note that a right R-module may be viewed as a left module over R^{∘}, and vice-versa.

An R-bimodule is an abelian group with compatible left and right R-module structures. One example of an R-bimodule is R itself, viewed as a left and right R-module via left and right multiplication respectively. In the sequel, this will be the canonical bimodule structure on R.

Suppose that R is commutative, that is to say, rs = (−1)^{|r||s|}sr for each r, s ∈ R. Then an R-bimodule M is said to be symmetric if the left and right structures determine each other, in the sense that, for r ∈ R and m ∈ M, one has rm = (−1)^{|r||m|}mr. For example, R is itself symmetric. Moreover, any left (or right) R-module can be naturally enriched to a symmetric R-bimodule.

Next we recall the basics of projective and injective resolutions.

1.2. Resolutions. Let R be a DG algebra and M a left R-module. A left R-module P is K-projective if Hom_{R} (P, −) preserves quasiisomorphisms. A K-projective resolution of M is a quasiisomorphism of left R-modules P → M, where P is Kprojective. Analogously, a left R-module I is K-injective if Hom_{R} (−, I) preserves quasiisomorphisms; a K-injective resolution of M is a quasiisomorphism of left R-modules M → I such that I is K-injective.

It is well known that every module has a K-projective and a K-injective resolution, and that they are unique up to quasiisomorphism; for example, see [5], or [16]. Thus one can construct the derived functor − ⊗_{R}^{L} − of the tensor product functor, as well as the derived functor RHom_{R} (−, −) of the homomorphism functor. Let L be a right R-module, and let M and N be left R-modules. Set

Tor_{i}^{R} (L, M) = H_{i}(L ⊗_{R}^{L} M) and Ext_{R}^{i} (M, N) = H_{−i}(RHom_{R} (M, N) ) .

When R is a ring, these correspond with the classically defined objects.

Let X be an R-bimodule. A biprojective resolution of X is a quasiisomorphism of R-bimodules P → X such that P is K-projective both as a left R-module and as a right R-module. An biinjective resolution of X is defined analogously.

It turns out that for the specific modules that interest us biinjective resolutions always exist, but biprojective resolutions may not. This problem explains our interest in the next result. First, some notation.

Henceforth, 𝕜 denotes a fixed commutative noetherian ring. A DG algebra over 𝕜 is a DG algebra R equipped with a morphism of DG algebras 𝕜 → R with the property that the image of 𝕜 lies in the centre of R. Here 𝕜 is viewed as a DG algebra (necessarily concentrated in degree 0) with trivial differential.

For each complex X of 𝕜-modules, X^{♮} denotes the underlying graded 𝕜-module. 1.3. Proposition. Let R be a DG algebra over 𝕜 and let X be an R-bimodule. Then X has a biprojective resolution under either one of the following conditions.

(a) R is commutative and X is symmetric.

(b) R_{i} = 0 for i ≪ 0 and the 𝕜-module R^{♮} is projective.

(c) 𝕜 is a field.

Proof. (a) Let P → X be a K-projective resolution of the left R-module X. Since R is commutative, P can be endowed with a structure of a symmetric R-bimodule, and with this structure, the homomorphism P → X is one of R-bimodules. Moreover, P is K-projective also when considered as a right R-module. Thus, P is a biprojective resolution of X.

(b) Since X is a bimodule, it can be viewed as a left module over the enveloping DG algebra R^{e} = R ⊗_{𝕜} R^{∘}. Let P be a K-projective resolution of X over R^{e}; in particular, P → X is a quasiisomorphism of R-bimodules. Under our hypothesis, R is a K-projective (left) 𝕜-module, so by base change we deduce that R^{e} is a K-projective left R-module. Thus, the adjunction isomorphism:

Hom_{R} (P, −) ≅ Hom_{R} (R^{e} ⊗_{R^{e}} P, −) ≅ Hom_{R^{e}} (P, Hom_{R} (R^{e}, −) )

yields that P is K-projective on the left.

An analogous argument establishes that P is K-projective also on the right.

(c) This proof in this case is akin to the one in (b). □

The next few paragraphs concern finiteness issues for DG algebras and modules. 1.4. Finiteness. Let R be a DG algebra and M a (left or right) R-module.

We say that M is bounded to the left (respectively, bounded to the right) if H_{i}(M) = 0 for i ≫ 0 (respectively, for i ≪ 0); it is said to be bounded if it is bounded both to the left and to the right. The R-module M is degreewise finite if for each integer i the H_{0}(R)-module H_{i}(M) is finitely generated; M is finite if it is degreewise finite and bounded. In other words, if the graded H_{0}(R)-module H(M) is finitely generated.

One often encounters the following problem: One begins with a pair of (degreewise) finite modules and would like to know whether their derived tensor product has the same property. This turns out to be true in many situations of interest; this is the content of the next two results.

A DG algebra R is said to be connective if H_{i}(R) = 0 for i < 0, and coconnective if H_{i}(R) = 0 for i > 0. The following facts are well known; one way to prove them is to use the Eilenberg-Moore spectral sequence that converges, under either sets of hypotheses, from Tor^{H(R)} (H(L), H(M)) to Tor^{R} (L, M) ; see, for example, [5, (1.10)] or [7, (1.3.2)]. 1.5. Let R be DG 𝕜-algebra that is connective, and degreewise finite over 𝕜. Let L be a right R-module and M a left R-module. If L and M are degreewise finite and bounded to the right, then the same holds for the complex of 𝕜-modules L⊗_{R}^{L} M. □ 1.6. Let k be a field and R a degreewise finite coconnective DG k-algebra such that H_{0}(R) = k and H_{−1}(R) = 0. Let L be a right R-module and M a left R-module. If L and M are degreewise finite and bounded to the left, then the same holds for the complex of k-vector spaces L ⊗_{R}^{L} M. □

Next we wish to introduce, following [11], dualizing DG modules. To this end, it is expedient to have on hand the following notion.

1.7. Let R be a DG algebra and X an R-bimodule. We say that a left R-module M is X-reflexive if the following biduality morphism is an isomorphism:

M → RHom_{R^{∘}} (RHom_{R} (M, X) , X) .

A right R-module N is X-reflexive if it is X-reflexive over R^{∘}; in other words, if the biduality morphism N → RHom_{R} (RHom_{R^{∘}} (N, X) , X) is an isomorphism.

1.8. Dualizing modules. Let R be a DG algebra. An R-bimodule D is dualizing (for R) if for any finite left R-module M and finite right R-module N, the following conditions are satisfied.

(1) D has a biprojective resolution and a biinjective resolution;

(2) The R-modules RHom_{R} (M, D) and RHom_{R^{∘}} (N, D) are both finite;

(3) The left R-modules M and D ⊗_{R}^{L} M are D-reflexive, as are the right Rmodules N and N ⊗_{R}^{L} D;

(4) R is D-reflexive as left R-module and as a right R-module.

The properties asked of a dualizing module are inspired by the ones enjoyed by dualizing complexes over commutative noetherian rings; confer Hartshorne [14, Chapter V] or P. Roberts [15, Chapter 2]. Over such rings, the following strengthening of (1.8.3), contained in [14, V.2.1], holds; this is useful in the sequel.

1.9. Let C be a symmetric dualizing complex for 𝕜. If a complex of 𝕜-modules V is degreewise finite, then V is C-reflexive. □

For the remainder of this section let R be a DG 𝕜-algebra.

In Section 2 we prove that in many cases one can obtain a dualizing module over R by coinducing one from 𝕜. This calls for a recap on this process.

1.10. Coinduction. Let R be a DG 𝕜-algebra and V a complex of 𝕜-modules.

The complex of 𝕜-modules RHom_{𝕜} (R, V) acquires a structure of an R-bimodule as follows: Let I be a K-injective resolution of V. Since (the image of) 𝕜 is central in R, the complex of 𝕜-modules Hom_{𝕜} (R, I) inherits a structure of an R-bimodule from the canonical one on R. This R-bimodule structure does not depend on the choice of the injective resolution: If J is another K-injective resolution of V, then there is a quasiisomorphism I → J; this induces the quasiisomorphism Hom_{𝕜} (R, I) → Hom_{𝕜} (R, J) , which is a morphism of R-bimodules.

There are a couple of simple observations that are worth recording:

(a) If R is commutative, then the R-bimodule RHom_{𝕜} (R, V) is symmetric.

(b) If R is K-projective when viewed as a complex of 𝕜-modules, then the Rbimodules RHom_{𝕜} (R, V) and Hom_{𝕜} (R, V) are quasiisomorphic.

Indeed, the R-bimodule structure on RHom_{𝕜} (R, V) is induced by the one on Hom_{𝕜} (R, I) , where I is a K-injective resolution of V. Thus, (a) holds because the canonical R-bimodule structure on R is symmetric. As to (b): The canonical map Hom_{𝕜} (R, V) → Hom_{𝕜} (R, I) is a quasiisomorphism since R is K-projective over 𝕜. An elementary calculation verifies that this homomorphism is one of R-bimodules.

1.11. Let V be a complex of 𝕜-modules, and M a left R-module. The derived category avatar of the classical Hom-Tensor adjointness yield a quasiisomorphism

\xymatrix{RHom_{R} (M, RHom_{𝕜} (R, V) ) \ar[r]^{≃} & RHom_{𝕜} (M, V)} .

Now, RHom_{𝕜} (R, V) is an R-bimodule, so its right R-structure induces a left Rstructure on RHom_{R} (M, RHom_{𝕜} (R, V) ) . On the other hand, since the image of 𝕜 is central in R, the left R-module structure of M carries over to RHom_{𝕜} (M, V) . The quasiisomorphism above is compatible with this additional data.

2. Dualizing DG modules

The gist of this section is that many of the naturally occurring DG algebras possess dualizing modules. We begin by describing the main results; their proofs are to be found at the end of the section. 2.1. Theorem. Let A be a commutative noetherian ring and C a dualizing complex for A. Let K be the Koszul complex on a finite set of elements in A. Then the K-bimodule Hom_{A} (K, C) is dualizing. 2.2. Let φ: A′ → A be a local homomorphism, so that A′ and A are local rings and φ maps the maximal ideal of A′ to the maximal ideal of A. Let k denote the residue field of A′.

One can construct a DG A′-algebra resolution A′[X] ⟶^{≃} k with the property that X is a positively graded set and A′[X] is a free commutative A′-algebra on X. Such a resolvent is unique up to quasiisomorphism of DG A′-algebras. The reader may refer to, for instance, [2, (2.1)] for details on these matters.

Set A[X] = A⊗_{A′} A′[X]; this is called the DG fibre of φ. It is immediate from the uniqueness of resolvents that the DG fibre is well defined up to quasiisomorphism of DG algebras. Since A′[X] is a free resolution of k, the k-algebras H(A[X]) and Tor^{A′} (A, k) are isomorphic. Theorem. Let φ: A′ → A be a local homomorphism such that the flat dimension of A over A′ is finite, and let C be a dualizing complex for A. Let A[X] be the DG fibre of φ. Then the A[X]-bimodule Hom_{A} (A[X], C) is dualizing. Remark. For commutative DG algebras, Apassov [1, §4] has considered a weaker notion of dualizing modules and has proved that DG fibres possess such dualizing modules; see [1, (4.3)]. That result is contained in the theorem above. 2.3. Let P be a complex over a commutative ring A and set E = Hom_{A} (P, P) . Composition defines a product structure on E making it a DG algebra; this is the endomorphism DG algebra of P. Since A is commutative, homothety induces a map A → E; this is a homomorphism of DG algebras with A lying in the centre of E. In other words, E is an A-algebra.

For the next result, recall that P is said to be perfect if P^{♮} is bounded and, for each integer i, the A-module P_{i} is finitely generated and projective. Theorem. Let A be a commutative noetherian ring and C a dualizing complex for A. Let P be a perfect complex of A-modules with H(P) ≠ 0 and E the endomorphism DG algebra Hom_{A} (P, P) . Then the E-bimodule Hom_{A} (E, C) is dualizing. 2.4. Recall our standing assumption that 𝕜 denotes a commutative noetherian ring.

Let G be a topological monoid and let C_{∗} (G; 𝕜) denote the singular chain complex of G with coefficients in 𝕜. It is well known that the product on G induces a structure of a DG 𝕜-algebra on C_{∗} (G; 𝕜); for example, see [7, pp. 28,88].

Theorem. Let C be a dualizing complex for 𝕜. If the homology of G with coefficients in 𝕜 is finitely generated, then the C_{∗} (G; 𝕜)-bimodule Hom_{𝕜} (C_{∗} (G; 𝕜) , C) is dualizing.

The preceding theorems are deduced from more general results described below. 2.5. Proposition. Let R be a DG 𝕜-algebra that is finite over 𝕜, and C a dualizing complex for 𝕜. Set D = RHom_{𝕜} (R, C) . Suppose that the following conditions hold.

(1) D has a biprojective resolution;

(2) For any finite left R-module M (respectively, finite right R-module N), the

𝕜-module D ⊗_{R}^{L} M (respectively, N ⊗_{R}^{L} D) is C-reflexive. Then D is a dualizing module for R.

Proof. We verify that D has the properties required of it by (1.8).

Condition (1): The desired biprojective resolution is provided by our hypothesis. As to the biinjective one: Let J be an injective resolution of C over 𝕜. Then, the R-bimodule I = Hom_{𝕜} (R, J) is K-injective both on the left and the right, as is easily verified using the adjunction isomorphism in (1.11).

A few remarks before we proceed with the verification of (1.8.2) and (1.8.3): By symmetry, it suffices to establish the assertions concerning left R-modules. Also,

(a) For any left R-module L and right R-module N, (1.11) yields that:

RHom_{R} (L, D) ⟶^{≃} RHom_{𝕜} (L, C) and RHom_{R^{∘}} (N, D) ⟶^{≃} RHom_{𝕜} (N, C) .

(b) Since R is finite over 𝕜, a module over R is finite if and only if it is finite when viewed as a complex of 𝕜-modules.

Condition (2): As C is a dualizing complex for 𝕜, condition (1.8) yields that for any finite left R-module M, the R-module RHom_{𝕜} (M, C) is finite as well. It remains to note that, by (a), the latter module is isomorphic to RHom_{R} (M, D) .

Condition (3): We claim that if a left R-module L is C-reflexive when viewed as a complex of 𝕜-modules, then it is D-reflexive. This is justified by the following commutative diagram where θ and γ are the respective biduality morphisms, and α and β are (induced by) the adjunction quasiisomorphisms; see (a) above.

\xymatrix{L \ar[r]^{θ}_{≃} \ar[d]_{γ} & RHom_{𝕜} (RHom_{𝕜} (L, C) , C) \ar[d]_{≃}^{α} \\ _{RHom_{R^{∘}} (RHom_{R} (L, D) , D)} \ar[r]^{≃}_{β} & _{RHom_{𝕜} (RHom_{R} (L, D) , C)}}

Now, if M is a finite left R-module, then, viewed as complexes of 𝕜-modules, both M and D ⊗_{R}^{L} M are C-reflexive - the former by (1.9) and the latter by assumption - so it follows from the preceding discussion that they are D-reflexive as well.

Condition (4): Since R is finite over 𝕜, this is the special case M = R of the already verified Condition (3).

This completes our proof that D is dualizing for R. □

The preceding proposition allows us to establish the existence of dualizing modules for diverse classes of DG algebras. 2.6. Proposition. Let R be a connective DG 𝕜-algebra that is finite over 𝕜, and C a dualizing complex for 𝕜. If either R_{i} = 0 for i ≪ 0 and the 𝕜-module R^{♮} is projective, or R is commutative, then the R-bimodule RHom_{𝕜} (R, C) is dualizing.

Proof. Set D = RHom_{𝕜} (R, C) . Our hypothesis place us in the situation treated by (2.5). So it suffices to verify that D satisfies conditions (1) and (2) of loc. cit.

Condition (1): When R is commutative, D is symmetric; see (1.10). Thus, under either sets of hypotheses, (1.3) provides us with a biprojective resolution.

Condition (2): The 𝕜-module R is finite and C is a dualizing complex over 𝕜, so D is finite when viewed as a complex of 𝕜-modules. In particular, D is finite over R. Therefore, R being connective, for any finite left R-module M, the 𝕜-module H(D ⊗_{R}^{L} M) is degreewise finite and bounded to the right; see (1.5). Now (1.9) yields that D ⊗_{R}^{L} M is C-reflexive, as desired.

An analogous argument settles the assertion concerning right R-modules. □

This completes our preparation for proving the results announced at the beginning of this section.

Proof of Theorem 2.1. The DG A-algebra K is finite, connective, and commutative. Thus, by (2.6), the K-bimodule RHom_{A} (K, C) is dualizing. It remains to note that since K^{♮} is a bounded complex of projective A-modules, the K-bimodules Hom_{A} (K, C) and RHom_{A} (K, C) are quasiisomorphic; see (1.10). □

Proof of Theorem 2.2. The DG fibre A[X] of φ is a commutative DG A-algebra with structure map the canonical inclusion A → A[X]; see (2.2). The homology of A[X] is Tor^{A′} (A, k) , so A[X] is connective; it is also finite because the flat dimension of A′ over A is finite. Thus, RHom_{A} (A[X], C) is a dualizing module for A[X], by (2.6). It remains to observe that RHom_{A} (A[X], C) ≃ Hom_{A} (A[X], C) as A[X]-bimodules, since the A-module A[X] is free. □

Proof of Theorem 2.3. Let D = RHom_{A} (E, C) . Since E is a bounded complex of projective A-modules, the R-bimodule Hom_{A} (E, C) is quasiisomorphic to D, by (1.10). So it suffices to prove that D is dualizing for E. To this end, we verify that the DG A-algebra E meets the criteria set out in (2.5).

To begin with, note that E is finite over A; this is by construction.

Condition (1): For each i, the A-module E_{i} is projective and E_{i} = 0 when i ≪ 0, so (1.3) yields that D has a biprojective resolution.

Condition (2): In the following sequence of isomorphisms of right E-modules, the one on the left holds because E is a bounded complex of projective A-modules, the one in the middle is by [13], and the last one is the canonical one.

RHom_{A} (E, C) ≃ C ⊗_{A}^{L} RHom_{A} (E, A) ≃ C ⊗_{A}^{L} Σ^{n}E = Σ^{n}C ⊗_{A}^{L} E .

Let M be a finite left E-module. The quasiisomorphism above implies that the left module D ⊗_{E}^{L} M is quasiisomorphic to (Σ^{n}C ⊗_{A}^{L} E) ⊗_{E}^{L} M, that is to say, to Σ^{n}C ⊗_{A}^{L} M. By (1.5), this last module is degreewise finite and bounded to the right, hence (1.9) yields that it is C-reflexive, as desired.

An analogous argument may be used to verify that N ⊗_{E}^{L} D is C-reflexive for any right R-module N. □

Proof of Theorem 2.4. Set R = C_{∗} (G; 𝕜). By construction, R is concentrated in non-negative degrees, so it is connective. Moreover, the homology of G is the homology of the chain complex of G so the hypothesis translates to: The 𝕜-module H(R) is finitely generated, that is to say R is finite over 𝕜. Finally, since the 𝕜-module R^{♮} is projective, (2.6) yields that RHom_{𝕜} (R, C) is dualizing for R; it remains to invoke (1.10) to obtain that Hom_{𝕜} (R, C) is quasiisomorphic to RHom_{𝕜} (R, C) , and hence, dualizing for R. □

3. Dualizing modules over local DG algebras

This section focuses on properties of dualizing modules over commutative local DG algebras. First, the relevant definition.

A commutative DG algebra R is said to be local if

(1) R is concentrated in non-negative degrees: R = {R_{i}}_{i ⩾ 0};

(2) R_{0} is noetherian;

(3) H_{0}(R) is local and H(R) is degreewise finite. The residue field of such a DG algebra R is the residue field of the local ring H_{0}(R). The residue field k has the structure of an R-module via the canonical surjections R → H_{0}(R) → k. Remark. Typically, when dealing with commutative rings it is tacitly assumed that dualizing complexes are symmetric. And indeed all the result in this section are proved under such a hypothesis. However, we have opted not to incorporate this as part of our definition of dualizing modules (over commutative DG algebras); cf. (1.8). This is keeping with line with the point of view adopted here that commutative DG algebras are specializations of general DG algebras.

The following existence theorem for dualizing modules is a DG algebra version of [14, (V.3.4)]; it differs from Theorem II stated in the introduction only in the form. It can be used, for example, to give quick proofs of theorems 2.1 and 2.2. 3.1. Theorem. Let R be a finite commutative local DG algebra with residue field k, and let D be a symmetric R-bimodule. The following conditions are equivalent.

(1) D is dualizing.

(2) D is finite and rank_{k} Ext_{R} (k, D) = 1.

The theorem in proved in (3.6).

When R has a dualizing module, the preceding result is contained in the next one. It is a generalization of the well known result that (symmetric) dualizing complexes over a commutative local ring are quasiisomorphic up to suspension [14, Chapter 5. §3]. It contains also Theorem III from the introduction. 3.2. Theorem. Let R be a finite commutative local DG algebra, and D, E symmetric R-bimodules. If D is dualizing, then the following conditions are equivalent.

(1) E ≃ Σ^{m}D^{r} for some integers m and r.

(2) E is finite, Ext_{R} (k, E) is concentrated in a single degree, and it has rank r. In particular, if E is dualizing, then it is quasiisomorphic to D, up to suspension.

The proof is given in (3.9).

The following result is a first step towards theorems (3.1) and (3.2); note that it does not require that R be finite. Numerically, it can be interpreted as stating that, up to a shift, the Bass (respectively, Betti) numbers of D equal the Betti (respectively, Bass) numbers of R. In fact, one can extend this result such that it deals with any finite R-module M and its dual RHom_{R} (M, D) . The purpose on hand does not call for that generality. 3.3. Proposition. Let R be a commutative local DG algebra with residue field k, and let D be a symmetric dualizing R-bimodule. Then there is an integer n such that (a) RHom_{R} (k, D) ≃ Σ^{n}k.

(b) RHom_{R} (D, k) ≃ Σ^{−n}RHom_{R} (k, R) .

Proof. The action of R on RHom_{R} (k, D) - which is apriori induced from D - coincides with the action induced from k. This is because R is commutative and D is symmetric. Thus, as an R-module, RHom_{R} (k, D) is quasiisomorphic to a complex of k-vector spaces with trivial differential:

RHom_{R} (k, D) ≃ ⨁_{i∈ℤ} Σ^{i}k^{b_{i}} .

Since k is a finite R-module, and D is dualizing, RHom_{R} (k, D) is finite. Therefore, b_{i} = 0 for all but a finitely many i, so that

RHom_{R} (RHom_{R} (k, D) , D) ≃ ⨁_{i∈ℤ} Σ^{−i}RHom_{R} (k, D)^{b_{i}}

≃ ⨁_{i,j∈ℤ} Σ^{j−i}k^{b_{i}b_{j}} .

The biduality morphism k → RHom_{R} (RHom_{R} (k, D) , D) is a quasiisomorphism since k is finite. Thus, the quasiisomorphism above entails: There exists an integer n such that b_{n} = 1, whilst b_{i} = 0 for i ≠ n. This is the result we seek.

(b) In the following sequence of quasiisomorphisms, the first is given by (a), the third holds because R is commutative, while the fourth is by the definition of dualizing modules.

RHom_{R} (D, k) ≃ RHom_{R} (D, Σ^{−n}RHom_{R} (k, D) )
≃ Σ^{−n}RHom_{R} (D, RHom_{R} (k, D) )
≃ Σ^{−n}RHom_{R} (k, RHom_{R} (D, D) )
≃ Σ^{−n}RHom_{R} (k, R) .

This is the desired quasiisomorphism. □

Our proof of Theorem 3.1 is built on the following special case. It is well known when, in addition, C is bounded to the left. Although this is all that is required in the sequel, it might be worthwhile to record the more general statement. 3.4. Proposition. Let (S, 𝔪, k) be a local ring and C a degreewise finite complex of symmetric S-bimodules. If rank_{k} Ext_{S} (k, C) = 1, then C is dualizing.

Proof. Since rank_{k} Ext_{S} (k, C) is finite, sup{i | Ext_{S}^{i} (k, C) ≠ 0} is finite. This last number is, by definition, the depth of the complex C - see [9, (2.3)]. Thus, as C is degreewise finite, [loc. cit., (2.5)] yields that C is bounded above. Now apply, for example, [14, (V.3.4)]. □

Here is the last piece of machinery required in the proof of Theorem 3.1. 3.5. Truncations. Let R be a commutative local DG algebra and M a finite R-module with H(M) ≠ 0. Set s = sup{i | H_{i}(M) ≠ 0}, and

τ(M): ⋯ → M_{s+2} ⟶^{∂_{s+2}} M_{s+1} ⟶^{∂_{s+1}} Ker(∂_{s}) → 0 .

Apriori, the inclusion ι: τ(M) ⊆ M is only one of complexes of abelian groups. However, a direct calculation establishes that τ(M) is stable under multiplication by elements in R. Therefore, it inherits an R-module structure from M, and ι is a homomorphism of R-modules. It extends to a triangle

(†)  τ(M) ⟶^{ι} M → C(ι) →

in the derived category of R-modules.

Since H_{s}(M) = Ker(∂_{s})/∂_{s+1}(M_{s+1}), there is a surjection π: τ(M) → Σ^{s}H_{s}(M); it is a degree 0 homomorphism of R-modules, with R acting on H_{s}(M) via H_{0}(R).

We require the following observations. Recall that the amplitude of M, denoted amp(M), is defined to be sup{i | H_{i}(M) ≠ 0} − inf{i | H_{i}(M) ≠ 0}; it is finite since we have assumed that M is finite.

(a) π is a quasiisomorphism. Moreover, if amp(M) = 0, then ι: τ(M) → M is also a quasiisomorphism, so M ≃ Σ^{s}H_{s}(M) as R-modules.

(b) amp(τ(M)) = 0 and amp(C(ι)) ≤ amp(M) − 1.

(c) If M is (degreewise) finite, then so is C(ι). Furthermore, for any R-modules D and X one has:

(d) If both RHom_{R} (τ(M), D) and RHom_{R} (C(ι), D) are (degreewise) finite, then so is RHom_{R} (M, D) .

(e) If both X ⊗_{R}^{L} τ(M) and X ⊗_{R}^{L} C(ι) are D-reflexive, then so is X ⊗_{R}^{L} M. Indeed, by construction, π is a quasiisomorphism, H_{s}(ι): H_{s}(τ(M)) → H_{s}(M)

is bijective and H_{n}(τ(M)) = 0 for n ≠ s. This explains (a), and the equality amp(τ(M)) = 0 in (b). From (a) and the homology long exact sequence of H_{0}(R)- modules engendered by (†) above

⋯ → H_{n}(τ(M)) → H_{n}(M) → H_{n}(C(ι)) → H_{n−1}(τ(M)) → ⋯,

one obtains that H_{n}(C(ι)) ≅ H_{n}(M) for n ≤ s − 1 and 0 otherwise. This yields amp(C(ι)) ≤ amp(M) − 1, and also (c). As to (d), it is immediate from the homology exact sequence associated to the triangle

→ RHom_{R} (C(ι), D) → RHom_{R} (M, D) → RHom_{R} (τ(M), D)

obtained by applying RHom_{R} (−, D) to (†). Finally, let F denote the biduality functor RHom_{R} (RHom_{R} (−, D) , D) associated to D. The triangle (†) and the naturality of the biduality morphism yields the ladder:

\xymatrix{X ⊗_{R}^{L} τ(M) \ar[r] \ar[d]^{≃} & X ⊗_{R}^{L} M \ar[r] \ar[d] & X ⊗_{R}^{L} C(ι) \ar[d]^{≃} \\ F (X ⊗_{R}^{L} τ(M)) \ar[r] & F (X ⊗_{R}^{L} M) \ar[r] & F (X ⊗_{R}^{L} C(ι))}

where the quasiisomorphism are due to X⊗_{R}^{L}τ(M) and X⊗_{R}^{L}C(ι) being D-reflexive. The associated homology exact ladder and the five lemma yield that the middle vertical arrow is a quasiisomorphism as well; that is to say, X ⊗_{R}^{L} M is D-reflexive. 3.6. Proof of Theorem 3.1.

(1) ⟹ (2). Since R is finite and D is dualizing, RHom_{R} (R, D) is finite; that is to say, D is finite. The remaining assertion is contained in Proposition 3.3.a.

(2) ⟹ (1) We verify that D satisfies the conditions set out in (1.8). The basic strategy is to reduce everything to questions concerning modules over the ring H_{0}(R). To facilitate its implementation, let S = H_{0}(R) and C = RHom_{R} (S, D) . Note that C has a natural structure of a symmetric S-bimodule; see (1.10).

Claim. C is dualizing for S.

Indeed, for any S-module M, viewed as an R-module via the surjection R → S, adjointness yields the quasiisomorphism of R-modules

(*)  \xymatrix{RHom_{S} (M, C) \ar[r]^{≃} & RHom_{R} (M, D) .}

In particular, RHom_{S} (k, C) ≃ RHom_{R} (k, D) , so that rank_{k} Ext_{S} (k, C) = 1. Moreover, since R is connective, and both S and D are finite, C is degreewise finite and bounded to the left. It remains to invoke Proposition 3.4.

Note that since D is symmetric one need worry only about left R-modules. Also, for any R-module M the left and right structures on M ⊗_{R}^{L} D and RHom_{R} (M, D) coincide. These remarks will be used without further ado.

Condition (1): Since R is commutative and D is symmetric, any K-projective, respectively, K-injective, resolution can be enriched to a biprojective, respectively, biinjective, one.

Condition (2): If M is a finite S-module, then so is RHom_{S} (M, C) , since C is dualizing for S. Thus, RHom_{R} (M, D) is finite, by (*). This entails the finiteness of RHom_{R} (M, D) also when M is a finite R-module.

Indeed, the required finiteness is tautological if H(M) = 0, so assume that H(M) is nontrivial. Now induce on the amplitude of M. Suppose that amp M = 1. Then M ≃ Σ^{s}H_{s}(M) for some integer s, by (3.5.a), so by passing to H_{s}(M) one may assume that M is a finite S-module; this case has been dealt with.

Let n be a positive integer such that the desired finiteness holds for any finite R-module with amplitude at most n. Let M be a finite R-module with amp(M) = n + 1. Consider the R-modules τ(M) and C(ι) constructed as in (3.5). Both these have amplitude ≤ n, by (3.5.b), and are finite, by (3.5.c). Thus, the induction hypothesis yields that RHom_{R} (τ(M), D) and RHom_{R} (C(ι), D) are both finite. Now (3.5.d) yields that RHom_{R} (M, D) is finite as well.

Thus, Condition (2) is verified.

Condition (3): We check: If an R-module M is finite, then D⊗_{R}^{L}M is D-reflexive. To begin with, suppose that M is an S-module, with R acting on it via S. Then

the action of R on D ⊗_{R}^{L} M induced by D coincides with the one induced by M, so D ⊗_{R}^{L} M is quasiisomorphic, as an R-module, to a (DG) S-module U. Although U may not be finite, it is degreewise finite and bounded to the right; see (1.5).

The biduality morphism θ: U → RHom_{R} (RHom_{R} (U, D) , D) fits in the following commutative square. In it γ is the biduality map associated to U when viewed as an S-module; by (1.9), it is a quasiisomorphism since U is degreewise finite and C is dualizing. Adjunction - see (*) - induces the remaining quasiisomorphisms.

\xymatrix{U \ar[r]^{θ} \ar[d]^{≃}_{γ} & RHom_{R} (RHom_{R} (U, D) , D) \ar[d]^{≃} \\ _{RHom_{S} (RHom_{S} (U, C) , C)} \ar[r]^{≃} & _{RHom_{R} (RHom_{S} (U, C) , D)}}

So θ is a quasiisomorphism; that is to say, U, and hence D ⊗_{R}^{L} M, is D-reflexive.

Now an induction argument akin to the one used in verifying Condition (2) settles the issue for any finite R-module M. Here is a sketch: If amp(M) = 0, then, by passing to a quasiisomorphic R-module, one may assume that M is an S-module; this case has already been settled. If amp(M) ≥ 1, then the R-modules τ(M) and C(ι) from (3.5.c) are both finite, and have amplitudes less than amp(M). So the induction hypothesis ensures that both D ⊗_{R}^{L} τ(M) and D ⊗_{R}^{L} C(ι) are D-reflexive. Therefore, (3.5.e) - with X = D - yields that D ⊗_{R}^{L} M is also D-reflexive.

A similar argument establishes that any finite R-module is D-reflexive.

Condition (4): This is covered by Condition (3) as R is finite.

Thus D is dualizing. This completes the proof of the theorem. □ Now to prepare for the proof of Theorem 3.2.

3.7. Let R be a commutative local DG algebra and M an R-module. If M is degreewise finite and bounded to the right, then one can construct a resolution F of M such that there is an isomorphism of R^{♮}-modules

F^{♮} ≅ ⨁_{i∈ℤ} Σ^{i}(R^{♮})^{b_{i}} with b_{i} = 0 for i ≪ 0 ,

and the complex F ⊗_{R} k has trivial differential. Such an F is said to be a minimal semifree resolution of M; see [4, §10], [7, (A.3)].

Here is a simple application of the preceding construction. 3.8. Lemma. Let R be a commutative local DG algebra with residue field k, and let M be an R-module that is degreewise finite and bounded to the right. If there integers n and b such that Ext_{R} (M, k) ≅ Σ^{n}k^{b}, then M ≃ Σ^{−n}R^{b}.

Proof. Let F be a minimal semifree resolution of M given by (3.7). Then the complex of k-vector spaces Hom_{R} (F, k) , being isomorphic to Hom_{k} (F ⊗_{R} k, k) , has trivial differential. The isomorphism Ext_{R} (M, k) ≅ Σ^{n}k^{b} translates to the isomorphism: Hom_{R} (F, k) ≅ Σ^{n}k^{b}. Therefore, b_{i} = 0 for i ≠ −n whilst b_{−n} = b. In other words, F ≅ Σ^{−n}R^{b}; thus M ≃ Σ^{−n}R^{b}. □

3.9. Proof of Theorem 3.2. (1) ⟹ (2): Since R is finite and D is dualizing, RHom_{R} (R, D) is finite; thus D is finite. Therefore, E is also finite. Moreover

RHom_{R} (k, E) ≃ RHom_{R} (k, Σ^{m}D^{r}) ≃ Σ^{m}RHom_{R} (k, D)^{r} ≃ Σ^{m+n}k^{r} ,

while the last quasiisomorphism is due to Proposition 3.3.a. Taking homology yields the desired conclusion.

(2) ⟹ (1): Since E is finite and D is dualizing, RHom_{R} (E, D) is finite. Moreover, the following biduality morphism is a quasiisomorphism.

(†)  E ⟶^{≃} RHom_{R} (RHom_{R} (E, D) , D) .

This engenders the first of the following sequence of quasiisomorphisms; the second is due to the commutativity of R, whilst the third is by (3.3).

RHom_{R} (k, E) ≃ RHom_{R} (k, RHom_{R} (RHom_{R} (E, D) , D) )
≃ RHom_{R} (RHom_{R} (E, D) , RHom_{R} (k, D) )
≃ RHom_{R} (RHom_{R} (E, D) , Σ^{n}k)
≃ Σ^{n}RHom_{R} (RHom_{R} (E, D) , k) .

The hypothesis is that H_{l}(RHom_{R} (k, E) ) ≅ k^{r} for some integers l, r. This is tantamount to: RHom_{R} (k, E) ≃ Σ^{l}k^{r}; see (3.5.a). Thus, the quasiisomorphism above yields that RHom_{R} (RHom_{R} (E, D) , k) ≃ Σ^{l−n}k^{r}. Now Lemma (3.8) implies that RHom_{R} (E, D) ≃ Σ^{n−l}R^{r}. Feeding this into equation (†) yields:

E ≃ RHom_{R} (Σ^{n−l}R^{r}, D) ≃ Σ^{l−n}D^{r} .

This completes the proof that (2) ⟹ (1).

Finally, if E is dualizing, then rank_{k} Ext_{R} (k, E) = 1, by Proposition 3.3.a, so the already established implication (2) ⟹ (1) yields that E ≃ Σ^{m}D, as desired. □

In attempting to extend Theorem 3.2 to dualizing modules that are not symmetric, one has to contend with the following phenomenon. Example. Let R be a commutative local ring and σ: R → R a nontrivial ring automorphism. Let R^{σ} denote the R-bimodule on the set underlying R with multiplication defined by r · x = rx and x · r = xσ(r), for r ∈ R and x ∈ R^{σ}. Then R^{σ} and R are not quasiisomorphic as bimodules.

Indeed, suppose there is such a quasiisomorphism. This induces a bimodule isomorphism φ: R^{σ} → R. Hence, for each r ∈ R one has

φ(r − σ(r)) = φ(r · 1 − 1 · r) = rφ(1) − φ(1)r = 0 .

In particular, r = σ(r); this contradicts the non-triviality of σ.

On the other hand, if R is Gorenstein, then both R and R^{σ} are dualizing; the first by definition, and the second by a direct verification.

We hope to investigate this matter elsewhere.

4. Gorenstein DG algebras

In this short section we apply the results in Section 3 to the study of Gorenstein DG algebras, as introduced in [13]. To begin with, a definition. 4.1. A DG algebra R is Gorenstein if R is dualizing, in sense of (1.8). Since R is biprojective and R-reflexive, it follows from the definition that R is Gorenstein if and only if for each finite left R-module M and finite right R-module N:

(1) R has a biinjective resolution;

(2) The R-modules RHom_{R} (M, R) and RHom_{R^{∘}} (N, R) are both finite;

(3) M and N are R-reflexive. 4.2. Remark. Let R be a commutative local DG algebra with residue field k, as defined in (3.3). If R is Gorenstein, then Ext_{R} (k, R) ≅ Σ^{n}k for some integer n; this is by (3.3.a).

In [3] Avramov and Foxby define a commutative local DG algebra R to be ‘Gorenstein’ if it is finite and the k-vector space Ext_{R} (k, R) has rank 1. Thus, among finite commutative local DG algebras, the Gorenstein class as identified by [13] is potentially smaller than that identified by [3]. In fact, these classes coincide; this is immediate from Theorem 3.1. We record this fact in the following result; it contains Theorem I from the introduction. 4.3. Theorem. Let R be a finite commutative local DG algebra with residue field k. The following conditions are equivalent.

(1) R is Gorenstein;

(2) rank_{k} Ext_{R} (k, R) = 1. □ As a corollary, we obtain the following result which completes [13, (2.11)].

4.4. Let Q be a commutative local DG algebra with residue field k, and let φ: Q → T be a homomorphism of DG algebras such that φ(Q) lies in the centre of T and T ⊗_{Q} k ≠ 0. Assume that the following conditions hold:

(1) T has a biinjective resolution over itself; (2) Viewed as an Q-module, T admits a K-projective resolution P ⟶^{≃} T such

that P ⊗_{Q} k has trivial differential. We refer the reader to [13] for the meanings of the terms ‘finite’ and ‘Gorenstein’.

Theorem. Suppose φ is finite and Gorenstein. Then, Q is Gorenstein if and only if T is Gorenstein.

Proof. Suppose Q is Gorenstein. Since Q is commutative, it has a biinjective resolution; this follows from an argument analogous to that of 1.3.a., and T has a biinjective resolution, by hypothesis. Thus [13, (2.6)] yields that T is Gorenstein.

If T is Gorenstein, then according to [13, (2.11)], the k-vector space Ext_{Q} (k, Q) is one dimensional. It remains to invoke Theorem 4.3. □

5. Cochain complexes

The results in this section are geared towards applications in topology. For this reason, all graded objects that appear will be graded cohomologically. For instance, any complex M is assumed to be of the form {M^{i}}_{i∈ℤ} with differential that increases degrees: ∂^{i}: M^{i} → M^{i+1}.

For the rest of this section, let k be a field. The following definitions are motivated by the example of cochain complexes of topological spaces; cf. (5.5). 5.1. A DG k-algebra R is said to be cochain if R is concentrated in non-negative degrees: R = {R^{i}}_{i ⩾ 0} ; such an R is simply connected if R^{0} = k and H^{1}(R) = 0.

Given a simply connected cochain DG algebra R, we call k its residue field. It has the structure of an R-module via the canonical surjection R → k.

Cochain DG algebras are cohomological versions of local DG algebras introduced in Section 3. With one major difference: they are not assumed to be commutative.

The following result ensures that cochain DG algebras posses dualizing modules. Its proof is akin to that of Proposition 2.6, except that, in verifying Condition 2 of (2.5) one invokes (1.6) instead of (1.5). Recall that R is said to be degreewise finite if for each integer i, the k-vector space H^{i}(R) is finite dimensional; it is finite if, in addition, H(R) is bounded. 5.2. Proposition. Let R be a finite simply connected cochain DG k-algebra. Then the DG R-bimodule Hom_{k} (R, k) is dualizing. □

The remaining results concern cochain DG k-algebras that are Gorenstein, that is to say, those that are dualizing; see (4.1). To begin with, one has the following counterpart of implication (2) ⟹ (1) of Theorem 4.3. 5.3. Theorem. Let R be a finite simply connected cochain DG k-algebra such that H(R) is commutative. If rank_{k} Ext_{R} (k, R) = 1, then R is Gorenstein.

Proof. Set D = Hom_{k} (R, k) ; by (5.2), this R-module is dualizing.

Claim. It suffices to prove that D is quasiisomorphic to a suspension of R as a left R-module and also as a right R-module.

Indeed, assume that D has this property. We verify conditions set out in (4.1). Condition (1): This follows from arguments analogous to the proof of (1.3.c). By symmetry, it is enough to check the remaining conditions for left R-modules. Condition (2): Say D ≃ Σ^{n}R as left R-modules, for some integer n. Then, for

any finite left R-module M, the k-modules RHom_{R} (M, D) and Σ^{n}RHom_{R} (M, R) are quasiisomorphic. Since D is dualizing, this yields that the latter module is finite over k, and hence over R, because H^{0}(R) = R^{0} = k.

Condition (3): Let M be a finite left R-module. The quasiisomorphism in the diagram below is by adjunction, while evaluation yields the first homomorphism.

RHom_{R} (M, R) ⟶^{ε} RHom_{R} (D ⊗_{R}^{L} M, D) ⟶^{≃} RHom_{R} (M, RHom_{R} (D, D) ) .

The composed homomorphism coincides with the one induced by the homothety R → RHom_{R} (D, D) . Since this last map is a quasiisomorphism one obtains that ε is also one such. Note that ε is compatible with the right R-module structures. This results in the second of the following quasiisomorphisms; the first one is the biduality morphism, and hence also a quasiisomorphism.

D ⊗_{R}^{L} M ⟶^{≃} RHom_{R^{∘}} (RHom_{R} (D ⊗_{R}^{L} M, D) , D)
⟶^{≃} RHom_{R^{∘}} (RHom_{R} (M, R) , D) .

Therefore, the composed map γ: D ⊗_{R}^{L} M → RHom_{R^{∘}} (RHom_{R} (M, R) , D) is a quasiisomorphism. Note that γ is the canonical evaluation homomorphism. Since D and R are quasiisomorphic as right R-modules and the left R-module structure of D is not involved in the definition of γ, we deduce that the corresponding evaluation morphism M = R ⊗_{R} M → RHom_{R^{∘}} (RHom_{R} (M, R) , R) , which is the biduality morphism, is a quasiisomorphism. Thus, Condition (3) is satisfied.

This completes the justification of our claim.

It remains to prove that D and R are quasiisomorphic, up to suspension, as left R-modules and also as right R-modules. To this end, note that since k is a field, H(D) = Hom_{k} (H(R), k) . By [7, (3.6)], our hypothesis entails that H(R) satisfies Poincaré duality, that is to say, the H(R)-module H(D) is isomorphic to Σ^{n}H(R), for some integer n. Observe that, since H(R) is commutative, H(D) is symmetric and the last isomorphism respects this structure.

The Eilenberg-Moore spectral sequence that converges from Tor^{H(R)} (k, H(D)) to Tor^{R} (k, D) - see [7, (1.3.2)] - collapses and the k-vector space Tor^{R} (k, D) is one dimensional. This entails D ≃ Σ^{n}R-module as left R-modules; this follows by an argument akin to that in the proof of Proposition 3.8.

Similarly, one deduces that D ≃ Σ^{n}R also as right R-modules.

This completes the proof of the theorem. □

Here is a partial converse to the preceding theorem; it is the cohomological version of (the special case D = R) of (3.3.a), and can be proved in the same way. 5.4. Theorem. Let R be a degreewise finite simply connected cochain DG k-algebra. Suppose that R is commutative. If R is Gorenstein, then rank_{k} Ext_{R} (k, R) = 1. □

Given Theorem 5.3, one is tempted to raise the following Question. In the result above, can one weaken the hypothesis ‘R is commutative’ to ‘H(R) is commutative’?

The last result is a topological avatars of the ones above. 5.5. Let X be a topological space. The complex of cochains on X with coefficients in k, which is denoted C^{∗} (X; k), is a DG k-algebra, with product defined by the Alexander-Whitney map; see [8, §5]. Although this product is not itself commutative, it is homotopy commutative, so that the cohomology of C^{∗} (X; k) is a commutative k-algebra. It is denoted H^{∗}(X; k).

According to Félix, Halperin, and Thomas a topological space X is Gorenstein at k if the k-vector space Ext_{C}_{∗}_{(X;k)} (k, C^{∗} (X; k)) is one dimensional.

5.6. Theorem. Let X be a simply connected topological space such that the k-vector space H^{∗}(X; k) is finite dimensional. If X is Gorenstein at k, then the DG algebra C^{∗} (X; k) is Gorenstein. The converse holds if the characteristic of k is 0.

Proof. Since H(C^{∗} (X; k)) is the cohomology of X, our hypotheses translate to: C^{∗} (X; k) is a finite, simply connected, cochain DG k-algebra. Thus, by Theorem 5.3, if X is Gorenstein at k, then C^{∗} (X; k) is Gorenstein.

Suppose that the characteristic of k is 0. Then C^{∗} (X; k) is quasiisomorphic to a commutative DG k-algebra R; see [7, §10]. It immediate from the definition that C^{∗} (X; k) and R are Gorenstein simultaneously; moreover, Ext_{C}_{∗}_{(X;k)} (k, C^{∗} (X; k)) and Ext_{R} (k, R) are isomorphic. The desired result is thus contained in (5.4). □

References

[1] D. Apassov, Dualizing modules over differential graded rings, pp. 41–47 in “Complexes and Differential Graded Modules”, ph.d. thesis, Lund University, 1999.

[2] L. L. Avramov, Infinite free resolutions, in: Six lectures in commutative algebra (Bellaterra, 1996), Progress in Math. 166, Birkhäuser, Boston, 1998; pp. 1–118.

[3] L. L. Avramov, H.-B. Foxby, Locally Gorenstein homomorphisms, Amer. J. Math. 114 (1992), 1007–1047.

[4] L. L. Avramov, H.-B. Foxby, S. Halperin, Differential graded homological algebra, in preparation.

[5] L. L. Avramov, S. Halperin, Through the looking glass: a dictionary between rational homotopy theory and local algebra, in: Algebra, algebraic topology and their interactions (proceedings of the conference in Stockholm, 1983), Lecture Notes in Math., Vol. 1183, Springer, Berlin, 1986; pp. 1–27.

[6] W. Dwyer, J. P. C. Greenlees, S. Iyengar, Duality in algebra and topology, preprint (2002).

[7] Y. Félix, S. Halperin, J.-C. Thomas, Gorenstein spaces, Adv. Math. 71 (1988), 92–112.

[8] Y. Félix, S. Halperin, J.-C. Thomas, “Rational Homotopy Theory”, Grad. Texts in Math., Vol. 205, Springer, Berlin, 2001.

[9] H.-B. Foxby, S. Iyengar, Depth and amplitude of unbounded complexes, in: Commutative Algebra. Its interaction with Algebraic Geometry (Grenoble-Lyon 2001), Contemporary Math., to appear.

[10] A. Frankild, P. Jørgensen, Homological identities for Differential Graded Algebras, II , preprint (2002).

[11] A. Frankild, P. Jørgensen, Dualizing DG-modules for Differential Graded Algebras, preprint (2001).

[12] A. Frankild, P. Jørgensen, Homological identities for Differential Graded Algebras, J. Algebra, to appear.

[13] A. Frankild, P. Jørgensen, Gorenstein Differential Graded Algebras, Israel J. Math., to appear.

[14] R. Hartshorne, “Residues and duality”, Lecture Notes in Math., Vol. 20, Springer, Berlin, 1966, Lecture notes of a seminar on the work of A. Grothendieck, given at Harvard 1963/64. With an appendix by P. Deligne.

[15] P. Roberts, Homological invariants of modules over commutative rings, Sèm. Math. Sup., Univ. Montréal, 1980.

[16] N. Spaltenstein, Resolutions of unbounded complexes, Compositio Math. 65 (1988), no. 2, 121–154.

Matematisk Afdeling, Københavns Universitet, København Ø, DK–Danmark

E-mail address: frankild@math.ku.dk

Mathematics Department, University of Missouri, Columbia, MO 65211, USA

E-mail address: iyengar@math.missouri.edu

Danish National Library of Science and Medicine, Nørre Allé 49, 2200 København N, DK–Danmark

E-mail address: pej@dnlb.dk